OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form. Answer y = 3x + 6 y = 3x + 3 y = 3x – 3 y = 3x – 6

OpenStudy (anonymous):

y=3x-3

OpenStudy (aroub):

-3-6/0-3 -9/-3 3 y=mx+b , m= 3 ,( 3,6) 6=3(3)+b 6=9+b -3=b so the equation is y= 3x-3

OpenStudy (anonymous):

@ aroub one medal plz